Brian is 4 years older than Michelle. Five years ago the sum of their ages was 48. How old are they now?
Found 2 solutions by Boreal, fractalier:
Answer by Boreal(15235)

Brian is x
Michelle is x-4
five years ago, Brian was x-5 and Michelle x-9.
Those two add to 48
x-5+x-9=48
2x-14=48
2x=62
x=31
Brian is 31, Michelle is 27 now
Five years ago, they were 26 and 22 respectively, and that sum is 48.

Answer by fractalier(6550)

Call their ages NOW, b and m.
Then we write
b = m + 4
Now five years ago, we had
(b-5) + (m-5) = 48
We can plug in for b and get
(m+4-5) + (m-5) = 48
(m-1) + (m-5) = 48
2m - 6 = 48
2m = 54
m = 27 years old
and then
b = 31 years old